### v1

# HarmoQ: Harmonized Post-Training Quantization for High-Fidelity Image Super-Resolution

Abstract
Post-training quantization offers an efficient pathway to deploy super-resolution models, yet existing methods treat weight and activation quantization independently, missing their critical interplay. Through controlled experiments on SwinIR, we uncover a striking asymmetry: weight quantization primarily degrades structural similarity, while activation quantization disproportionately affects pixel-level accuracy. This stems from their distinct roles—weights encode learned restoration priors for textures and edges, whereas activations carry input-specific intensity information. Building on this insight, we propose HarmoQ, a unified framework that harmonizes quantization across components through three synergistic steps: structural residual calibration proactively adjusts weights to compensate for activation-induced detail loss, harmonized scale optimization analytically balances quantization difficulty via closed-form solutions, and adaptive boundary refinement iteratively maintains this balance during optimization. Experiments show HarmoQ achieves substantial gains under aggressive compression, outperforming prior art by 0.46 dB on Set5 at 2-bit while delivering 3.2× speedup and 4× memory reduction on A100 GPUs. This work provides the first systematic analysis of weight-activation coupling in super-resolution quantization and establishes a principled solution for efficient high-quality image restoration.
Introduction
Single-image super-resolution (SISR) has witnessed remarkable progress driven by deep neural networks, especially Transformer-based architectures such as SwinIR (Liang et al. 2021) and Restormer (Zamir et al. 2022). These models achieve impressive reconstruction accuracy but come at a steep cost: their high parameter counts and computational demands pose major barriers to deployment in real-world scenarios, particularly in edge or mobile environments.
Post-training quantization (PTQ) offers an appealing solution to compress pretrained models without retraining. However, applying PTQ to super-resolution remains an open challenge. Prior works diverge on how to balance quantization: some target both weights and activations (Liu et al. 2024), while others focus exclusively on activations (Wang et al. 2025). Yet, the relative impact of these two components on visual fidelity remains poorly understood.
Figure 1: Weight vs. activation quantization analysis. (a) Performance comparison showing that W32A4 (activation quantization) primarily degrades PSNR while W4A32 (weight quantization) shows stronger SSIM degradation, indicating complementary effects on pixel-level accuracy versus structural similarity. (b, c) Visual comparison of W32A4 and W4A32 quantization results on anime image reconstruction.
To systematically investigate this trade-off, we conduct a controlled comparison between two complementary 4-bit quantization regimes—W4A32 (quantized weights, full-precision activations) and W32A4 (full-precision weights, quantized activations)—on SwinIR across standard benchmarks. As shown in Figure 1, a clear pattern emerges: weight quantization (W4A32) has a stronger effect on SSIM (drops to 0.85 on Manga109), indicating corruption of fine-grained textures and edge structures, while activation quantization (W32A4) disproportionately affects PSNR (drops to 32.5 dB on Urban100), reflecting degradation of pixel-level accuracy and global attributes such as brightness and contrast.
Key Observation: Weight quantization primarily impacts structural similarity and texture fidelity (reflected in SSIM degradation). Activation quantization mainly affects pixel-level accuracy and global appearance attributes (reflected in PSNR loss).
This asymmetry arises from the distinct functional roles of weights and activations. Weights encode distributed, global restoration priors learned during training; their quantization introduces systematic errors that primarily corrupt representations of fine textures and edges. In contrast, activations carry local, input-specific intensity information; their quantization introduces stochastic noise that disproportionately affects pixel-level accuracy, degrading global attributes like brightness uniformity and overall contrast.
Figure 2: Layer-wise quantization sensitivity analysis in SwinIR. Stacked bar chart shows the proportion of sensitivity to weight versus activation quantization across four layer types. Attention and GELU layers are highly sensitive to activation quantization (92.5% and 93.6%), while Shallow layers show balanced sensitivity. Attention maps compare (a) W32A4 and (b) W4A32 configurations, illustrating differential quantization impacts in attention layer.
Despite these findings, most existing quantization methods either treat weight and activation quantization independently, as in DBDC (Tu et al. 2023) and 2DQuant (Liu et al. 2024), or focus on a single component, as in dynamic activation-only quantization (Wang et al. 2025). These methods overlook the intrinsic coupling between weights and activations. Classical PTQ techniques such as MinMax (Jacob et al. 2018) or Percentile (Li et al. 2019) further exacerbate this issue by applying uniform clipping or scaling heuristics that ignore signal content related to structure and texture, leading to suboptimal quantization behavior.
To address these challenges, we introduce HarmoQ, a harmonized post-training quantization framework designed specifically for super-resolution models. HarmoQ jointly minimizes the compound error caused by weight-activation interaction by incorporating structure-aware mechanisms. First, it introduces a structural residual calibration scheme that adjusts weights to proactively compensate for structural degradation arising from activation quantization. Second, it derives a closed-form optimal scaling factor that analytically balances the quantization difficulty across components. Third, it applies adaptive boundary refinement to maintain this balance consistently throughout optimization, even under aggressive quantization regimes.
This work contributes the first systematic analysis of the asymmetric effects of weight versus activation quantization in super-resolution. It also proposes a closed-form structural calibration technique to mitigate the loss of fine-grained details, and an analytically grounded scaling strategy for harmonizing quantization across components. Together, these advances lead to strong empirical gains in low-bit regimes, offering a robust and theoretically motivated solution for efficient super-resolution deployment.
Related Work
Image Super-resolution.
Image super-resolution aims to reconstruct high-resolution images from low-resolution inputs, with applications in medical imaging (Greenspan 2008; Isaac and Kulkarni 2015) and surveillance (Zhang et al. 2010). Early CNN approaches like SRCNN (Dong et al. 2014) pioneered deep learning for super-resolution. EDSR (Lim et al. 2017) achieved state-of-the-art results using enhanced residual networks, while SRGAN (Ledig et al. 2017) introduced adversarial training for photo-realistic results. RCAN (Zhang et al. 2018a) and RDN (Zhang et al. 2018b) further improved performance through attention mechanisms and dense connections. Recent transformer-based methods have shown promising results. SwinIR (Liang et al. 2021) adapted Swin Transformer (Liu et al. 2021) for image restoration, while Restormer (Zamir et al. 2022) proposed efficient transformer architectures for high-resolution restoration. CAT (Chen et al. 2022) and DAT (Chen et al. 2023b) explore cross-aggregation and dual-aggregation mechanisms for enhanced feature learning.
Model Quantization.
Model quantization reduces computational complexity by representing weights and activations with lower precision, enabling efficient deployment on resource-constrained devices. Early quantization-aware training methods include BinaryNets (Courbariaux et al. 2016), which constrained weights to binary values, and DoReFa-Net (Zhou et al. 2016), which extended low-bitwidth quantization. PACT (Choi et al. 2018) introduced parameterized clipping for quantized networks. Post-training quantization offers practical advantages by quantizing pre-trained models without retraining. BRECQ (Li and et al. 2021) achieved significant improvements through block-wise reconstruction, while recent work (Ding et al. 2022) addresses PTQ challenges for vision transformers. For super-resolution specifically, PAMS (Li et al. 2020) proposed parameterized max scale quantization, DAQ (Hong et al. 2022b) developed channel-wise distribution-aware methods, and CADYQ (Hong et al. 2022a) introduced content-aware dynamic quantization. Recent advances (Tu et al. 2023) have made progress toward accurate post-training quantization for super-resolution models.
Problem Formulation
Consider a linear transformation in a super-resolution network: $y=Wx+b,$ where $W\in\mathbb{R}^{m\times n}$ and $x\in\mathbb{R}^{n}$ denote the weights and input activation, respectively. Let $\tilde{W}$ and $\tilde{x}$ be their quantized counterparts, with corresponding quantization errors $\delta_{W}=\tilde{W}-W$ and $\delta_{x}=\tilde{x}-x$. The quantized forward pass expands to:
$\tilde{y}=Wx+W\delta_{x}+\delta_{W}x+\delta_{W}\delta_{x}+b.$ (1)
Our analysis in Figure 1 reveals that weight and activation quantization exhibit complementary degradation patterns: weight quantization primarily impacts structural similarity (degrading SSIM), while activation quantization corrupts pixel-level accuracy (degrading PSNR). This metric differentiation makes independent optimization suboptimal, as existing methods (Jacob et al. 2018; Tu et al. 2023) fail to minimize the critical compound error term $W\delta_{x}+\delta_{W}x$.
We propose a harmonized quantization framework that jointly optimizes quantization parameters to preserve critical structural and pixel-level information. Our approach formulates the problem as:
$\min_{s}$ $\mathcal{L}_{\mathrm{total}}(s)=\mathbb{E}\left[\left\|W\delta_{x}(s)+\delta_{W}(s)x\right\|_{2}^{2}\right],$ (2)
s.t. $\left|\mathrm{MSE}_{x}(s)-\mathrm{MSE}_{w}(s)\right|\leq\epsilon,$
where $s$ is a scaling parameter shared across quantization modules, and $\epsilon$ enforces distortion symmetry. The optimal $s^{*}$ satisfies $\mathrm{MSE}_{x}(s^{*})\approx\mathrm{MSE}_{w}(s^{*})$, yielding minimal total error under balanced quantization.
Figure 3: Structural Residual Calibration analysis in SwinIR. (a) Optimal weight calibration $\Delta W^{*}$ heatmap using Laplacian filter, revealing structured channel relationships. (b) Response comparison before/after $\Delta W^{*}$ calibration showing structure-aware modulation. (c) Efficacy across layers for different projection matrices. Laplacian filter achieves the best performance (see Table 3 for quantitative comparison), while random projection fails due to lack of structure-aware design. Cost analysis via Frobenius norm of $\Delta W$ shows consistent magnitude across projection types.
Figure 4: Comparison of quantization optimization strategies. (a) Existing methods (2DQuant, DOBI) independently optimize weight and activation quantization boundaries through separate calibration processes, leading to suboptimal parameter selection. (b) Our HarmoQ framework employs a unified three-step optimization: calibration for initial range estimation, harmonized scaling using optimal factor $s^{*}$ to balance quantization difficulty, and iterative moving bounds refinement to jointly minimize compound quantization errors. The arrows indicate the iterative optimization flow between scaling and boundary adjustment steps.
Methodology
As illustrated in Figure 4, traditional approaches like 2DQuant and DOBI (Figure 4a) calibrate weight and activation quantization parameters separately, resulting in suboptimal boundary choices that neglect their interplay. In contrast, our HarmoQ framework (Figure 4b) introduces a unified three-stage optimization strategy: (1) calibration to set initial parameter ranges, (2) harmonized scaling to balance quantization difficulty via an optimal scale factor $s^{*}$, and (3) adaptive boundary refinement to jointly adjust clipping ranges, minimizing the compound error. This coordinated process preserves critical image fidelity components and achieves more robust super-resolution performance.
Step 1: Structural Residual Calibration
In quantized super-resolution networks, the compound error term $W\delta_{x}+\delta_{W}x$ exhibits metric-specific degradation patterns. The core challenge stems from weight quantization errors $\delta_{W}$, which systematically corrupt the learned structural priors (e.g., for edges and textures) in $W$. This degradation is particularly severe in SR tasks where preserving edge sharpness and texture details is critical. Unlike weight quantization that introduces systematic biases affecting global image properties uniformly, activation quantization errors vary both spatially and temporally, making them particularly detrimental to the preservation of fine-grained visual details that are critical for super-resolution quality.
We propose a Structural Residual Calibration (SRC) mechanism that adjusts weight parameters to compensate for structural degradation introduced by weight quantization $\delta_{W}$ itself. This approach leverages the insight that weight parameters, with their distributed representation capacity, can be strategically modified to encode compensatory information that mitigates activation-induced loss of fine-grained detail before they manifest in the output.
Subspace projection operator.
Let $H\in\mathbb{R}^{k\times d}$ be a linear projection matrix that extracts structural features like edges and textures from $d$-dimensional output features. This matrix can be constructed using fixed filters such as DCT high-pass masks or Laplacian edge filters bases derived from calibration data.
Structure-aware calibration objective.
We seek a calibration term $\delta_{W}$ that suppresses the projected residual energy while maintaining stability through regularization:
$\min_{\delta_{W}}\mathbb{E}\left[\|H(W\delta_{x}+\delta_{W}x)\|_{2}^{2}\right]+\lambda\|\delta_{W}\|_{F}^{2},$ (3)
where $\lambda>0$ balances calibration strength and smoothness.
Closed-form solution.
By expanding the objective and taking its derivative with respect to $\delta_{W}$, we obtain the following optimality condition:
$\delta_{W}\left(H\mathbb{E}[xx^{T}]H^{T}+\lambda I\right)=-W\mathbb{E}[\delta_{x}x^{T}]H^{T}.$ (4)
Solving for $\delta_{W}$ yields the closed-form structure-aware calibration:
$\delta_{W}^{*}=-W\mathbb{E}[\delta_{x}x^{T}]H^{T}\left(H\mathbb{E}[xx^{T}]H^{T}+\lambda I\right)^{-1}$ (5)
The complete derivation of this result is provided in Appendix.
Projection analysis.
To validate our structural residual calibration, we analyze the optimal weight calibration $\Delta W^{*}$ across different projection matrices. Figure 3(a) shows the heatmap of $\Delta W^{*}$ derived via a Laplacian filter, revealing structured, non-random patterns that target specific channel relationships, confirming the structure-specific corrective capability of our method. Figure 3(b) illustrates how $\Delta W^{*}$ modulates activation responses while preserving key signal characteristics. The smooth transitions in the corrected profile support our structure-aware calibration principle and demonstrate reduced structural distortion. Figure 3(c) further shows consistent error reduction across layers for various structural projection filters. DCT high-pass and learned basis filters yield superior results, especially in deeper layers where quantization errors accumulate, while random projections perform poorly, underscoring the importance of a structure-aware design.
Step 2: Harmonized Scale Optimization
Following the structural calibration, we address the fundamental challenge of non-uniform quantization difficulty across weights and activations due to their distinct statistical properties. We seek a harmonizing scale factor $s$ that balances quantization difficulty between components.
Scale harmonization objective.
Given clipping boundaries $\theta=\{\alpha_{x},\beta_{x},\alpha_{w},\beta_{w}\}$, we seek a scale factor $s$ that ensures equal quantization difficulty: $\mathrm{MSE}_{x}(s)=\mathrm{MSE}_{w}(s).$ In SR networks, the activation range is further modulated by an input scaling factor $s$, which affects MSE differently for activations and weights:
$\mathrm{MSE}_{x}(s)=\frac{(\beta_{x}-\alpha_{x})^{2}}{12s^{2}(2^{b_{x}}-1)^{2}},\ \mathrm{MSE}_{w}(s)=\frac{(\beta_{w}-\alpha_{w})^{2}s^{2}}{12(2^{b_{w}}-1)^{2}}.$
Setting $\mathrm{MSE}_{x}(s)=\mathrm{MSE}_{w}(s)$ and solving for $s$ gives the closed-form optimal scale:
$s^{*}=\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}.$ (6)
The detailed derivation of the optimal scale factor is given in Appendix. This balances the difficulty of quantizing weights and activations, particularly in structure-sensitive regimes, ensuring that neither component dominates the quantization error.
Step 3: Adaptive Boundary Refinement
With the harmonized scale $s^{*}$ established, we optimize the quantization boundaries $\theta$ to minimize the total reconstruction error while maintaining the balanced quantization difficulty achieved in Step 2.
Boundary optimization objective.
We minimize the total quantization error by updating the quantization boundaries $\theta$:
$\mathcal{L}_{\mathrm{total}}(s^{*},\theta)=\mathbb{E}\left[\left\|W\cdot\delta_{x}(s^{*},\theta)+\delta_{W}(s^{*},\theta)\cdot x\right\|_{2}^{2}\right].$
The gradient with respect to any boundary parameter $\theta_{i}\in\theta$ is:
$\frac{\partial\mathcal{L}_{\mathrm{total}}}{\partial\theta_{i}}=\mathbb{E}\left[2\left(W\delta_{x}+\delta_{W}x\right)^{T}\cdot\left(W\frac{\partial\delta_{x}}{\partial\theta_{i}}+\frac{\partial\delta_{W}}{\partial\theta_{i}}x\right)\right].$
Gradient computation for clipping boundaries.
Since $\delta_{x}$ and $\delta_{W}$ are functions of the quantizer’s clipping range, their derivatives w.r.t. $\alpha$ and $\beta$ depend on the quantization function. For a symmetric uniform quantizer $Q(z)=\mathrm{clip}(z;\alpha,\beta)$ with step size $\Delta=\frac{\beta-\alpha}{2^{b}-1}$, we can compute:
$\frac{\partial\delta}{\partial\alpha}=\frac{\partial}{\partial\alpha}(Q(z)-z),\qquad\frac{\partial\delta}{\partial\beta}=\frac{\partial}{\partial\beta}(Q(z)-z),$
which are differentiable almost everywhere except at clipping points. The complete gradient derivation for boundary optimization is presented in Appendix.
Boundary update rule.
The boundaries are updated using projected gradient descent:
$\theta^{(t+1)}=\mathrm{Proj}_{\Omega}\left(\theta^{(t)}-\eta\nabla_{\theta}\mathcal{L}_{\mathrm{total}}(s^{*},\theta^{(t)})\right),$ (7)
where $\Omega$ is a feasible set ensuring $\alpha<\beta$ and $\eta$ is the learning rate.
Iterative Refinement Process
The unified framework alternates between three coordinated operations:
  1. 1.
    Step 1 - Structural Residual Calibration: Applies closed-form structural calibration $\delta W_{\ell}^{*}=-W_{\ell}\Sigma_{\delta x}H^{T}(H\Sigma_{xx}H^{T}+\lambda I)^{-1}$ to minimize structural distortion.
  2. 2.
    Step 2 - Harmonized Scale Optimization: Computes optimal scale factor $s^{*}=\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}$ to maintain equal quantization difficulty.
  3. 3.
    Step 3 - Adaptive Boundary Refinement: Updates clipping boundaries $\theta$ via gradient descent while enforcing the constraint $|\text{MSE}_{x}(s^{*},\theta)-\text{MSE}_{w}(s^{*},\theta)|\leq\epsilon$.
Table 1: Quantitative comparison of HarmoQ against SOTA quantization methods on standard SR benchmarks.
Set5 ($\times 2$) Set14 ($\times 2$) BSD100 ($\times 2$) Urban100 ($\times 2$) Set5 ($\times 4$) Set14 ($\times 4$) BSD100 ($\times 4$) Urban100 ($\times 4$)
Method Bit PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
SwinIR-light 32 38.15/0.9611 33.86/0.9206 32.31/0.9012 32.76/0.9340 32.45/0.8976 28.77/0.7858 27.69/0.7406 26.48/0.7980
Bicubic 32 32.25/0.9118 29.25/0.8406 28.68/0.8104 25.96/0.8088 27.56/0.7896 25.51/0.6820 25.54/0.6466 22.68/0.6352
MinMax 2 33.88/0.9185 30.81/0.8748 29.99/0.8535 27.48/0.8501 23.96/0.4950 22.92/0.4407 22.70/0.3943 21.16/0.4053
Percentile 2 30.82/0.8016 28.80/0.7616 27.95/0.7232 26.30/0.7378 23.03/0.4772 22.12/0.4059 21.83/0.3816 20.45/0.3951
DBDC+Pac 2 34.55/0.9386 31.12/0.8912 30.27/0.8706 27.63/0.8649 25.01/0.5554 23.82/0.4995 23.64/0.4544 21.84/0.4631
DOBI 2 35.25/0.9361 31.72/0.8917 30.62/0.8699 28.52/0.8727 28.82/0.7699 26.46/0.6804 25.97/0.6319 23.67/0.6407
Granular-DQ 2 35.85/0.9485 31.85/0.8995 30.80/0.8795 28.45/0.8800 29.35/0.8355 26.75/0.7305 26.35/0.6910 23.70/0.6895
2DQuant 2 36.00/0.9497 31.98/0.9012 30.91/0.8810 28.62/0.8819 29.53/0.8372 26.86/0.7322 26.46/0.6927 23.84/0.6912
HarmoQ 2 36.46/0.9515 32.24/0.9037 31.12/0.8836 29.18/0.8893 30.23/0.8543 27.29/0.7446 26.70/0.7031 24.27/0.7123
MinMax 3 28.19/0.6961 26.40/0.6478 25.83/0.6225 25.19/0.6773 19.41/0.3385 18.35/0.2549 18.79/0.2434 17.88/0.2825
Percentile 3 34.37/0.9170 31.04/0.8646 29.82/0.8339 28.25/0.8417 27.55/0.7270 25.15/0.6043 24.45/0.5333 22.80/0.5833
DBDC+Pac 3 35.07/0.9350 31.52/0.8873 30.47/0.8665 28.44/0.8709 27.91/0.7250 25.86/0.6451 25.65/0.6239 23.45/0.6249
DOBI 3 36.37/0.9496 32.33/0.9041 31.12/0.8836 29.65/0.8967 29.59/0.8237 26.87/0.7156 26.24/0.6735 24.17/0.6880
Granular-DQ 3 37.20/0.9555 32.75/0.9095 31.50/0.8900 30.30/0.9070 30.75/0.8690 27.65/0.7560 26.90/0.7115 24.70/0.7340
2DQuant 3 37.32/0.9567 32.85/0.9106 31.60/0.8911 30.45/0.9086 30.90/0.8704 27.75/0.7571 26.99/0.7126 24.85/0.7355
HarmoQ 3 37.98/0.9602 34.04/0.9220 32.23/0.9004 33.01/0.9368 31.09/0.8740 27.91/0.7618 27.11/0.7186 25.31/0.7548
HAT-S 32 38.58/0.9628 34.70/0.9261 32.59/0.9050 34.31/0.9459 32.92/0.9047 29.15/0.7958 27.97/0.7505 27.87/0.8346
Bicubic 32 32.25/0.9118 29.25/0.8406 28.68/0.8104 25.96/0.8088 27.56/0.7896 25.51/0.6820 25.54/0.6466 22.68/0.6352
MinMax 2 34.63/0.9263 31.23/0.8798 30.59/0.8595 30.95/0.8777 25.95/0.5746 24.16/0.4888 23.78/0.4483 22.37/0.4734
Percentile 2 32.06/0.8756 29.89/0.8118 28.97/0.7842 28.68/0.8288 24.18/0.5398 22.89/0.4507 22.34/0.4162 21.01/0.4346
DBDC+Pac 2 35.89/0.9457 32.29/0.9001 31.28/0.8777 30.61/0.8934 26.97/0.6285 25.16/0.5493 24.56/0.5044 23.18/0.5127
DOBI 2 36.63/0.9413 32.93/0.9008 31.69/0.8769 31.68/0.9004 31.04/0.8297 27.51/0.7104 26.89/0.6859 25.13/0.7089
Granular-DQ 2 36.70/0.9525 32.55/0.9070 31.30/0.8870 30.05/0.9045 30.30/0.8495 27.30/0.7405 26.80/0.7005 25.00/0.7410
2DQuant 2 36.81/0.9538 32.66/0.9085 31.42/0.8883 30.21/0.9061 30.48/0.8513 27.42/0.7423 26.89/0.7019 25.14/0.7425
HarmoQ 2 38.34/0.9617 34.35/0.9246 32.14/0.8961 33.87/0.9434 31.18/0.8749 27.95/0.7626 27.12/0.7191 25.33/0.7560
MinMax 3 30.86/0.7707 27.61/0.6957 26.07/0.6607 27.44/0.7567 21.77/0.4146 20.02/0.3287 19.77/0.3061 19.96/0.3764
Percentile 3 35.66/0.9283 32.17/0.8789 30.42/0.8453 31.37/0.8938 29.36/0.7855 26.13/0.6542 25.19/0.5998 25.16/0.7056
DBDC+Pac 3 36.41/0.9436 32.63/0.8977 30.99/0.8719 31.70/0.9018 29.82/0.7951 26.55/0.6769 25.71/0.6264 25.47/0.7178
DOBI 3 37.77/0.9558 33.49/0.9156 31.70/0.8901 33.09/0.9274 31.38/0.8586 27.83/0.7466 26.94/0.7085 26.89/0.8002
Granular-DQ 3 37.00/0.9520 33.15/0.9115 31.40/0.8860 32.65/0.9225 30.90/0.8520 27.45/0.7405 26.60/0.7030 26.40/0.7935
2DQuant 3 37.11/0.9532 33.26/0.9128 31.52/0.8872 32.78/0.9239 31.02/0.8534 27.57/0.7421 26.71/0.7043 26.54/0.7948
HarmoQ 3 38.35/0.9618 34.42/0.9250 32.36/0.9021 34.21/0.9441 31.90/0.8821 28.34/0.7700 27.41/0.7268 26.71/0.8106
Experiments
Baseline Methods.
We compare HarmoQ against several categories of quantization methods. Traditional approaches include MinMax (Jacob et al. 2018) and Percentile (Li et al. 2019) quantization, which use activation extrema and percentiles respectively to determine quantization boundaries. General post-training quantization methods include BRECQ (Li and et al. 2021), which pushes quantization limits through block-wise reconstruction, and DOBI, a data-free method that independently optimizes weight and activation quantization. For super-resolution specific methods, we evaluate against PAMS (Li et al. 2020) (parameterized max scale quantization), DBDC+Pac (Tu et al. 2023) (accurate post-training quantization for image super-resolution), 2DQuant (Liu et al. 2024) (low-bit post-training quantization), and Dynamic Granularity (Wang et al. 2025) (activation-only quantization strategy).
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Patch SwinIR-Light (3bit) MinMax 2DQuant DBDC
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference Image LR Patch DOBI Granular-DQ HarmoQ (w.o. SRC) HarmoQ (Ours)
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
GT Patch HAT-S (3bit) MinMax 2DQuant DBDC
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Reference Image LR Patch DOBI Granular-DQ HarmoQ (w.o. SRC) HarmoQ (Ours)
Figure 5: Visual comparison of different quantization methods on benchmark dataset.
Table 2: Component-wise ablation analysis.
Components Set5 ($\times 2$) Urban100 ($\times 4$)
SRC HSO ABR Config PSNR SSIM PSNR SSIM
SwinIR-Light 35.12 0.9387 23.12 0.6234
SRC 36.24 0.9468 23.89 0.6567
HSO 36.45 0.9482 24.15 0.6634
ABR 35.89 0.9434 23.56 0.6423
SRC+HSO 37.12 0.9531 24.67 0.7012
SRC+ABR 36.78 0.9503 24.23 0.6789
HSO+ABR 37.01 0.9521 24.45 0.6934
Full HarmoQ 37.98 0.9602 25.31 0.7548
Experimental Setup.
We evaluate HarmoQ on standard super-resolution benchmarks using SwinIR-light (Liang et al. 2021) and HAT-S (Chen et al. 2023a) as backbone networks. Following common practice, we train on DF2K dataset (combination of DIV2K (Timofte et al. 2017) and Flickr2K (Lim et al. 2017)) and test on five widely-used benchmarks: Set5 (Bevilacqua et al. 2012), Set14 (Zeyde et al. 2010), B100 (Martin et al. 2001), Urban100 (Huang et al. 2015), and Manga109 (Matsui et al. 2017). We conduct experiments with ×2 and ×4 upscaling factors under 2-bit and 3-bit quantization settings. Performance is measured using PSNR and SSIM computed on the Y channel of YCbCr color space.
Implementation Details.
We use Adam optimizer with learning rate 1×10-2, batch size 32, and 3000 total iterations. The structural projection matrix $H$ is constructed using Laplacian filters (see Table 3 for ablation), which empirically outperform DCT high-pass and learned bases by 0.26 dB on Set5. The regularization parameter is set to $\lambda=0.01$. Convergence is achieved when $|\mathcal{L}_{\mathrm{total}}^{(t+1)}-\mathcal{L}_{\mathrm{total}}^{(t)}|<10^{-4}$, typically within 15-20 iterations.
Quantitative Evaluation.
Table 1 presents comprehensive comparisons of HarmoQ against existing quantization methods on standard super-resolution benchmarks. Our method achieves consistent improvements under 2-bit and 3-bit quantization, particularly on structure-rich datasets like Urban100. This validates our core hypothesis that weight quantization affects structural similarity (SSIM) while activation quantization corrupts pixel-level accuracy (PSNR). Through unified optimization of structural residual calibration, harmonized scale balancing, and adaptive boundary refinement, HarmoQ transforms compound quantization errors from independent linear superposition to coordinated control. Cross-architecture experiments from SwinIR-light to HAT-S confirm the method’s generalizability, while performance on challenging 4× super-resolution demonstrates robustness under extreme compression. The quantitative results validate both HarmoQ’s technical superiority and the effectiveness of coupling analysis, providing a practical pathway for efficient super-resolution model deployment.
Table 3: Structural feature projection matrix analysis.
Set5 ($\times 2$) Urban100 ($\times 4$)
Projection Matrix $H$ PSNR SSIM PSNR SSIM
w/o SRC 37.01 0.9521 24.45 0.6934
Laplacian Filter 37.98 0.9602 25.31 0.7548
Sobel Edge Filter 37.85 0.9589 25.18 0.7521
DCT High-pass 37.72 0.9575 25.02 0.7495
Learned Basis 37.91 0.9594 25.24 0.7535
Random Projection 37.34 0.9548 24.78 0.7312
Component-wise ablation.
Our ablation study in Table 2 demonstrates that HarmoQ’s performance gains arise from complex nonlinear synergies rather than linear superposition. The baseline method, which processes weight and activation quantization in isolation while ignoring their coupling, embodies the dilemma of traditional super-resolution quantization. In contrast, our components show significant advantages. The SRC mechanism markedly improves performance by compensating for loss of structural information from activation quantization. HSO outperforms SRC when applied alone, highlighting the critical role of balancing quantization difficulty to enhance system robustness. ABR shows limited standalone effectiveness, as its performance depends on an established quantization balance. Pairwise combinations reveal strong synergy between SRC and HSO, where calibration is amplified by the quantization balance. Without HSO as a bridge, however, SRC and ABR struggle to synergize. Ultimately, the full HarmoQ framework outperforms all two-component combinations, demonstrating that its unified optimization enables a qualitative leap from local to global coordination. Consistent results on the Urban100 dataset validate this universality in scenes with rich textures.
Structural feature projection analysis.
Our ablation study in Table 2 explores the quest for optimal structure-aware calibration. Laplacian filters emerge victorious (37.98/0.9602), their second-order nature perfectly capturing edge discontinuities that activation quantization disrupts. Sobel filters (37.85/0.9589) and learned basis (37.91/0.9594) engage in a close contest, proving both classical gradient wisdom and data-driven intelligence can decode quantization artifacts. DCT high-pass filters (37.72/0.9575) show promise yet fall short, suggesting spatial gradients trump pure frequency decomposition. Random projection’s failure (37.34/0.9548) delivers the final verdict: effective error compensation demands explicit structural semantics, not blind dimensionality reduction.
Visual comparison.
Figure 5 demonstrates HarmoQ’s superiority under aggressive 3-bit quantization. MinMax produces severe artifacts with blurred giraffe fur and loss of building edge sharpness. DBDC and 2DQuant show improved structure but still exhibit texture degradation in fine details. Dynamic Granularity maintains better overall quality yet fails to preserve crisp window frames and fur texture boundaries. The ablation comparison reveals SRC’s critical contribution: HarmoQ (w.o. SRC) preserves the giraffe’s body contour but loses individual fur strand definition and produces softened building corner edges. Full HarmoQ recovers these fine-grained details—sharp fur textures, clear window boundaries, and preserved architectural edges—demonstrating that structural residual calibration effectively compensates for activation quantization artifacts in texture-sensitive regions.
Conclusion
This work addresses the deployment challenge of high-performance image SR models under aggressive compression. We proposed HarmoQ, a harmonized post-training quantization framework that unifies three strategies: structural residual calibration, harmonized scale optimization, and adaptive boundary refinement. HarmoQ enables practical deployment of high-quality SR models on resource-constrained devices, representing a significant advance in model compression for image restoration tasks.
Acknowledgments
This work was partially supported by JST-Mirai Program JPMJMI23G1, JSPS KAKENHI Grant Numbers 24KK0209, 24K22318, and 22H00529, the SPRING GX project of the University of Tokyo (grant number JPMJSP2108), the Research Institute of Trustworthy Autonomous Systems at Southern University of Science and Technology, the Jilin Provincial International Cooperation Key Laboratory for Super Smart City, and the Jilin Provincial Key Laboratory of Intelligent Policing.
Appendix A Appendix
Derivation of Structural Residual Calibration
Let $x\!\in\!\mathbb{R}^{d}$ be the input activation, $W\!\in\!\mathbb{R}^{m\times d}$ the original weight matrix, and $\delta_{x},\delta_{W}$ the quantization errors for activations and weights, respectively. A linear layer after quantization produces the perturbed response
$\tilde{y}\;=\;W(x+\delta_{x})+(\delta_{W})x\;=\;Wx+W\delta_{x}+\delta_{W}x.$ (1)
Objective. We wish to choose $\delta_{W}$ so that the structural residual $H\!\bigl(W\delta_{x}+\delta_{W}x\bigr)$ is minimized in mean–square sense, while keeping the weight perturbation small. Formally, we solve
$\min_{\delta_{W}}\;\mathcal{L}_{\mathrm{HF}}\;=\;\mathbb{E}\!\left[\lVert H(W\delta_{x}+\delta_{W}x)\rVert_{2}^{2}\right]+\lambda\lVert\delta_{W}\rVert_{F}^{2},$ (2)
where $H\!\in\!\mathbb{R}^{k\times d}$ projects onto a chosen subspace representing structural features and $\lambda\!>\!0$ balances distortion and regularization.
Second-order statistics.
Define the covariance and cross-covariance
$\Sigma_{xx}\;=\;\mathbb{E}[xx^{\top}],\quad\Sigma_{\delta x}\;=\;\mathbb{E}[\delta_{x}x^{\top}],\quad\Sigma_{\delta\delta}\;=\;\mathbb{E}[\delta_{x}\delta_{x}^{\top}].$ (3)
Loss expansion.
Using $\lVert A\rVert_{2}^{2}=\mathrm{tr}(AA^{\top})$ and the linearity of trace,
$\mathcal{L}_{\mathrm{HF}}$ $=\mathrm{tr}\!\Bigl(H\bigl(W\Sigma_{\delta\delta}W^{\top}+W\Sigma_{\delta x}\delta_{W}^{\top}\bigr)H^{\top}\Bigr)$
$\quad+\mathrm{tr}\!\Bigl(H\bigl(\delta_{W}\Sigma_{\delta x}^{\top}W^{\top}+\delta_{W}\Sigma_{xx}\delta_{W}^{\top}\bigr)H^{\top}\Bigr)+\lambda\lVert\delta_{W}\rVert_{F}^{2}.$
Matrix calculus identities $\partial\,\mathrm{tr}(A\delta_{W}^{\top})/\partial\delta_{W}=A$ and $\partial\,\mathrm{tr}(\delta_{W}B\delta_{W}^{\top})/\partial\delta_{W}=2\delta_{W}B$ give
$\frac{\partial\mathcal{L}_{\mathrm{HF}}}{\partial\delta_{W}}=2H^{\top}HW\Sigma_{\delta x}+2H^{\top}H\delta_{W}\Sigma_{xx}+2\lambda\delta_{W}.$ (4)
Normal equation.
Setting the gradient of (2) to zero yields
$H^{\top}H\delta_{W}\Sigma_{xx}+\lambda\,\delta_{W}=-\,H^{\top}HW\Sigma_{\delta x}.$ (5)
Closed-form solution.
Right-multiplying by $\Sigma_{xx}^{-1}$ and pre-multiplying by $(H^{\top}H)^{-1}$ (assuming $H$ has full row rank in its subspace),
$\delta_{W}=-\,W\Sigma_{\delta x}H^{\top}\!\bigl(H\Sigma_{xx}H^{\top}+\lambda I_{k}\bigr)^{-1}.$ (6)
This is the unique minimizer because (2) is strictly convex in $\delta_{W}$ for $\lambda>0$.
Approximation justification.
In deriving the closed-form solution, we neglect the second-order term $\delta_{W}\delta_{x}$ in the forward pass because: (1) under low-bit quantization, $\mathbb{E}[\|\delta_{W}\|_{F}]\ll\|W\|_{F}$ and $\mathbb{E}[\|\delta_{x}\|_{2}]\ll\|x\|_{2}$, making their product asymptotically negligible; (2) empirical validation on our calibration set shows this approximation introduces <0.01<0.01 dB error in final PSNR; (3) including this term would require solving a bilinear optimization problem without a closed-form solution, defeating the efficiency goal of PTQ.
Theorem 1 (Optimal Structural Correction).
[…]
Proof.
Equating the gradient of the objective to zero produces the linear system whose unique solution is given in the closed form, completing the proof. ∎
Derivation of Optimal Scale Factor
Theorem 2 (Harmonized Scale Optimization).
For uniform quantization with bit-widths $b_{x}$ and $b_{w}$, and clipping ranges $\alpha_{x},\beta_{x}]$ and $\alpha_{w},\beta_{w}]$, the optimal scale factor that equalizes quantization difficulty is:
$s^{*}=\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}$ (7)
Proof.
For uniform quantization with $b$ bits over range $\alpha,\beta]$, the quantization step size is:
$\Delta=\frac{\beta-\alpha}{2^{b}-1}$ (8)
The mean squared quantization error for uniformly distributed inputs is:
$\text{MSE}=\frac{\Delta^{2}}{12}=\frac{(\beta-\alpha)^{2}}{12(2^{b}-1)^{2}}$ (9)
In super-resolution networks, input scaling by factor $s$ affects activation and weight quantization differently: - Activation values are scaled by $s$, reducing their effective range to $\alpha_{x}/s,\beta_{x}/s] -$ Weight gradients are scaled by $s$, effectively expanding their range to $s\alpha_{w},s\beta_{w}]$
Therefore:
$\text{MSE}_{x}(s)$ $=\frac{(\beta_{x}/s-\alpha_{x}/s)^{2}}{12(2^{b_{x}}-1)^{2}}=\frac{(\beta_{x}-\alpha_{x})^{2}}{12s^{2}(2^{b_{x}}-1)^{2}}$ (10)
$\text{MSE}_{w}(s)$ $=\frac{(s\beta_{w}-s\alpha_{w})^{2}}{12(2^{b_{w}}-1)^{2}}=\frac{s^{2}(\beta_{w}-\alpha_{w})^{2}}{12(2^{b_{w}}-1)^{2}}$ (11)
Setting $\text{MSE}_{x}(s)=\text{MSE}_{w}(s)$:
$\frac{(\beta_{x}-\alpha_{x})^{2}}{12s^{2}(2^{b_{x}}-1)^{2}}=\frac{s^{2}(\beta_{w}-\alpha_{w})^{2}}{12(2^{b_{w}}-1)^{2}}$ (12)
Simplifying:
$\frac{(\beta_{x}-\alpha_{x})^{2}}{s^{2}(2^{b_{x}}-1)^{2}}=\frac{s^{2}(\beta_{w}-\alpha_{w})^{2}}{(2^{b_{w}}-1)^{2}}$ (13)
Cross-multiplying:
$(\beta_{x}-\alpha_{x})^{2}(2^{b_{w}}-1)^{2}=s^{4}(\beta_{w}-\alpha_{w})^{2}(2^{b_{x}}-1)^{2}$ (14)
Solving for $s$:
$s^{4}=\frac{(\beta_{x}-\alpha_{x})^{2}(2^{b_{w}}-1)^{2}}{(\beta_{w}-\alpha_{w})^{2}(2^{b_{x}}-1)^{2}}$ (15)
Taking the square root:
$s^{*}=\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}$ (16)
Gradient Computation for Boundary Optimization
Theorem 3 (Boundary Optimization Gradients).
For the total quantization error
$\mathcal{L}_{\text{total}}(s^{*},\theta)=\mathbb{E}\left[\left\|W\delta_{x}+\delta_{W}x\right\|_{2}^{2}\right],$
the gradient with respect to the clipping boundary $\theta_{i}$ is
$\frac{\partial\mathcal{L}_{\text{total}}}{\partial\theta_{i}}=2\,\mathbb{E}\left[(W\delta_{x}+\delta_{W}x)^{\!T}\left(W\frac{\partial\delta_{x}}{\partial\theta_{i}}+\frac{\partial\delta_{W}}{\partial\theta_{i}}x\right)\right].$
Proof.
The total quantization error can be written as
$\mathcal{L}_{\text{total}}=\mathbb{E}\left[\left(W\delta_{x}+\delta_{W}x\right)^{\!T}\left(W\delta_{x}+\delta_{W}x\right)\right].$
Taking the derivative with respect to $\theta_{i}$ gives
$\frac{\partial\mathcal{L}_{\text{total}}}{\partial\theta_{i}}$ $=\mathbb{E}\left[\frac{\partial}{\partial\theta_{i}}\left((W\delta_{x}+\delta_{W}x)^{\!T}(W\delta_{x}+\delta_{W}x)\right)\right]$
$=\mathbb{E}\left[\frac{\partial\left(W\delta_{x}+\delta_{W}x\right)^{\!T}}{\partial\theta_{i}}\left(W\delta_{x}+\delta_{W}x\right)\right]$
$+\mathbb{E}\left[\left(W\delta_{x}+\delta_{W}x\right)^{\!T}\frac{\partial\left(W\delta_{x}+\delta_{W}x\right)}{\partial\theta_{i}}\right].$
Since the inner product is symmetric,
$2\,\mathbb{E}\left[\left(W\delta_{x}+\delta_{W}x\right)^{\!T}\frac{\partial\left(W\delta_{x}+\delta_{W}x\right)}{\partial\theta_{i}}\right].$
Expanding the derivative inside,
$\frac{\partial\left(W\delta_{x}+\delta_{W}x\right)}{\partial\theta_{i}}=W\frac{\partial\delta_{x}}{\partial\theta_{i}}+\frac{\partial\delta_{W}}{\partial\theta_{i}}x+\delta_{W}\frac{\partial x}{\partial\theta_{i}}.$
Since $x$ does not depend on the quantization boundary,
$\frac{\partial x}{\partial\theta_{i}}=0,$
we finally have
$\frac{\partial\mathcal{L}_{\text{total}}}{\partial\theta_{i}}=2\,\mathbb{E}\left[\left(W\delta_{x}+\delta_{W}x\right)^{\!T}\left(W\frac{\partial\delta_{x}}{\partial\theta_{i}}+\frac{\partial\delta_{W}}{\partial\theta_{i}}x\right)\right].$
Convergence Analysis
Theorem 4 (Convergence of HarmoQ Algorithm).
Under mild regularity conditions, Algorithm 1 converges to a local minimum of the harmonized quantization objective.
Proof Sketch.
The proof follows from three key observations:
1. Monotonic Decrease: Each step of the algorithm decreases the objective function: - Step 1 (SRC) minimizes the structural residual energy - Step 2 (Scale Optimization) provides the closed-form optimal scale - Step 3 (Boundary Refinement) uses gradient descent with projection
2. Bounded Objective: The quantization error is bounded below by zero and above by the error of random quantization.
3. Lipschitz Gradients: Under standard assumptions on the data distribution, the gradients are Lipschitz continuous.
By the monotone convergence theorem and compactness of the feasible region, the algorithm converges to a stationary point of the constrained optimization problem. ∎
Complexity Analysis
Theorem 5 (Computational Complexity).
The per-iteration complexity of Algorithm 1 is $\mathcal{O}(kd^{2}+nd+L)$, where $k$ is the dimension of the structural feature subspace, $d$ is the feature dimension, $n$ is the number of parameters, and $L$ is the number of layers.
Proof.
The complexity analysis for each step:
Step 1 (SRC): - Computing $\mathbb{E}[xx^{T}]$: $\mathcal{O}(d^{2})$ - Computing $\mathbb{E}[\delta_{x}x^{T}]$: $\mathcal{O}(d^{2})$ - Matrix inversion $(H\Sigma_{xx}H^{T}+\lambda I)^{-1}$: $\mathcal{O}(k^{3})$ - Matrix multiplications: $\mathcal{O}(kd^{2})$
Step 2 (Scale Optimization): - Computing boundary statistics: $\mathcal{O}(n)$ - Scale factor computation: $\mathcal{O}(1)$
Step 3 (Boundary Refinement): - Gradient computation: $\mathcal{O}(nd)$ - Boundary updates: $\mathcal{O}(L)$
Total per-iteration complexity: $\mathcal{O}(kd^{2}+nd+L)$
Since typically $k\ll d$ for sparse structural feature projections, the dominant term is $\mathcal{O}(kd^{2})$. ∎
Algorithm 1 Harmonized Post-Training Quantization
0: Pre-trained model $\mathcal{M}$, calibration set $\mathcal{D}$, bit-widths $(b_{w},b_{x})$, tolerance $\tau$
0: Quantized model $\tilde{\mathcal{M}}$
1:Initialize: $\theta\leftarrow\text{MinMax}(\mathcal{D})$, $t\leftarrow 0$, $\mathcal{L}_{\text{prev}}\leftarrow\infty$
2:while $|\mathcal{L}_{\text{total}}^{(t)}-\mathcal{L}_{\text{prev}}|>\tau$ do
3:  $\mathcal{L}_{\text{prev}}\leftarrow\mathcal{L}_{\text{total}}^{(t)}$, $t\leftarrow t+1$
4:  Step 1: Structural Residual Calibration
5:  for each layer $\ell$ do
6:   $\Sigma_{xx}\leftarrow\mathbb{E}_{\mathcal{D}}[x_{\ell}x_{\ell}^{T}]$, $\Sigma_{\delta x}\leftarrow\mathbb{E}_{\mathcal{D}}[\delta_{x_{\ell}}x_{\ell}^{T}]$
7:   $\delta W_{\ell}^{*}\leftarrow-W_{\ell}\Sigma_{\delta x}H^{T}(H\Sigma_{xx}H^{T}+\lambda I)^{-1}$
8:   $W_{\ell}\leftarrow W_{\ell}+\delta W_{\ell}^{*}$
9:  end for
10:  Step 2: Harmonized Scale Optimization
11:  $s^{*}\leftarrow\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}$
12:  Step 3: Adaptive Boundary Refinement
13:  $\mathcal{L}_{\text{total}}\leftarrow\mathbb{E}_{\mathcal{D}}[\|W\delta_{x}(s^{*},\theta)+\delta_{W}(s^{*},\theta)x\|_{2}^{2}]$
14:  $\nabla_{\theta}\leftarrow\frac{\partial\mathcal{L}_{\text{total}}}{\partial\theta}$
15:  $\theta\leftarrow\text{Proj}_{\Omega}(\theta-\eta\nabla_{\theta})$ {s.t. $\alpha<\beta$}
16:  Constraint Enforcement:
17:  if $|\text{MSE}_{x}(s^{*},\theta)-\text{MSE}_{w}(s^{*},\theta)|>\epsilon$ then
18:   Recompute $s^{*}$ and re-apply Steps 1-2
19:  end if
20:end while
21:return Quantized model $\tilde{\mathcal{M}}$ with parameters $(s^{*},\theta^{*})$
Algorithm Details
The HarmoQ framework implements a unified three-stage iterative optimization strategy to address the coupling problem between weight and activation quantization in image super-resolution networks. As detailed in Algorithm 1, the algorithm begins by initializing quantization boundaries $\theta$ using MinMax calibration, then iteratively optimizes until convergence is achieved.
The algorithm operates through three coordinated steps. In Step 1 (lines 5-9), structural residual calibration computes the optimal weight calibration $\delta W_{\ell}^{*}=-W_{\ell}\Sigma_{\delta x}H^{T}(H\Sigma_{xx}H^{T}+\lambda I)^{-1}$ for each layer, where $H$ represents the structural feature projection matrix and $\lambda$ is the regularization parameter. This step preemptively compensates for the loss of structural information caused by activation quantization.
Step 2 (line 11) performs harmonized scale optimization by computing the optimal scale factor $s^{*}=\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}$ that ensures equal quantization difficulty between weights and activations, satisfying $\mathrm{MSE}_{x}(s^{*})=\mathrm{MSE}_{w}(s^{*})$.
Step 3 (lines 13-15) implements adaptive boundary refinement through projected gradient descent, minimizing the total quantization error $\mathcal{L}_{\mathrm{total}}=\mathbb{E}[\|W\delta_{x}(s^{*},\theta)+\delta_{W}(s^{*},\theta)x\|_{2}^{2}]$ while maintaining feasibility constraints.
The dynamic coupling mechanism (lines 17-19) ensures sustained coordination: when boundary updates violate the quantization balance ($|\mathrm{MSE}_{x}(s^{*},\theta)-\mathrm{MSE}_{w}(s^{*},\theta)|>\epsilon$), the algorithm automatically recomputes the optimal scale factor and reapplies the calibration steps. This creates a mutually reinforcing optimization trajectory where all three components remain harmonized throughout the process.
Algorithm 1 converges when the objective function change falls below threshold $\tau$, typically within 10-20 iterations. The per-iteration complexity is $\mathcal{O}(kd^{2}+nd+L)$, where the dominant term $\mathcal{O}(kd^{2})$ comes from the matrix operations in the structural calibration step.
Table 1: Detailed implementation specifications and hyperparameters.
Quantization Configuration
Weight quantization Per-channel symmetric, INT2/INT3/INT8
Activation quantization Per-tensor symmetric, INT2/INT3/INT8
Clipping scheme Learnable boundaries $\alpha,\beta]$ initialized by MinMax
Rounding mode Round-to-nearest-even (banker’s rounding)
Gradient estimator Straight-Through Estimator (STE)
Zero-point Fixed at 0 (symmetric quantization)
Quantization formula $Q(x)=\text{clip}\left(\text{round}\left(\frac{x}{s}\right),-2^{b-1},2^{b-1}-1\right)\cdot s$
Calibration Setup
Calibration set 1000 images from DIV2K training set
Image resolution $\times128$ RGB patches (HR)
Selection strategy Random sampling without replacement
Batch size 32 (limited by GPU memory)
Total iterations 3000 (typically converges at $\sim$2000)
Early stopping Enabled if $|\Delta\mathcal{L}|<10^{-5}$ for 50 consecutive iterations
Forward passes per batch 5 (for statistical estimation)
SRC Parameters
Projection matrix $H$ Laplacian filter (default), $\times3$ kernel: $\begin{bmatrix}0&1&0\\ 1&-4&1\\ 0&1&0\end{bmatrix}$
Alternative filters Sobel, DCT high-pass ($\times8$), learned basis
Subspace dimension $k$ 64 (shallow layers), 128 (deep layers), adaptive per block
Regularization $\lambda$ $1\times 10^{-2}$ (weight calibration), $5\times 10^{-3}$ (activation)
Covariance estimation Running average with momentum 0.9, 200 warmup samples
Matrix inversion Cholesky decomposition with $\epsilon=10^{-6}$ numerical stabilizer
Scale Optimization
Initial scale $s^{(0)}$ 1.0 (identity scaling)
Scale bounds $[0.1,10.0]$ to prevent numerical instability
Recomputation frequency Every 5 boundary update iterations
Balance tolerance $\epsilon$ $0.01\times\text{MSE}_{\text{initial}}$ (1% of initial error)
Closed-form solution $s^{*}=\sqrt{\frac{(\beta_{x}-\alpha_{x})(2^{b_{w}}-1)}{(\beta_{w}-\alpha_{w})(2^{b_{x}}-1)}}$
Boundary Refinement
Optimizer Adam with $\beta_{1}=0.9$, $\beta_{2}=0.999$, $\epsilon=10^{-8}$
Learning rate $\eta$ $1\times 10^{-2}$ (initial), cosine decay to $1\times 10^{-4}$
Learning rate schedule Cosine annealing over 3000 iterations, 300 warmup steps
Weight decay $1\times 10^{-5}$
Gradient clipping Global norm clipping at 1.0 to prevent divergence
Convergence threshold $\tau$ $1\times 10^{-4}$ (relative change in $\mathcal{L}_{\text{total}}$)
Feasibility projection $\alpha\leftarrow\min(\alpha,\beta-0.01)$, $\beta\leftarrow\max(\beta,\alpha+0.01)$
Iterative Coordination
Maximum iterations 3000 (with early stopping)
Rebalancing period $T$ Every 5 boundary update iterations
SRC reapplication Triggered when $|\text{MSE}_{x}-\text{MSE}_{w}|>1.5\epsilon$
Constraint violation handling Roll back to previous valid state, reduce $\eta$ by 0.5
Hardware and Environment
GPU NVIDIA A100 (40GB), CUDA 11.8, cuDNN 8.6
Framework PyTorch 2.0.1 with custom CUDA kernels for quantization
Precision Mixed precision (FP16 for forward, FP32 for gradient accumulation)
Memory optimization Gradient checkpointing enabled for large models
Calibration time 12.3 minutes (SwinIR-light), 18.7 minutes (HAT-S) on A100
Inference latency 14.1ms per image ($\times512$) for INT8 quantized model

### v2

@@
 This work provides the first systematic analysis of weight-activation coupling in super-resolution quantization and establishes a principled solution for efficient high-quality image restoration.
+Codehttps://github.com/Dreamzz5/HarmoQ
 Introduction
@@
 Figure 3: Structural Residual Calibration analysis in SwinIR.
-(a) Optimal weight calibration $\Delta W^{*}$ heatmap using Laplacian filter, revealing structured channel relationships.
-(b) Response comparison before/after $\Delta W^{*}$ calibration showing structure-aware modulation.
-(c) Efficacy across layers for different projection matrices.
-Laplacian filter achieves the best performance (see Table 3 for quantitative comparison), while random projection fails due to lack of structure-aware design.
+The top-left panel presents the optimal weight calibration $\Delta W^{*}$ heatmap using Laplacian filter, revealing structured channel relationships.
+The top-right panel shows response comparison before and after $\Delta W^{*}$ calibration, demonstrating structure-aware modulation.
+The bottom panel illustrates the efficacy across layers for different projection matrices.
+Laplacian filter achieves superior performance (see Table 3 for quantitative comparison), while random projection fails due to lack of structure-aware design.
 Cost analysis via Frobenius norm of $\Delta W$ shows consistent magnitude across projection types.
@@
 To validate our structural residual calibration, we analyze the optimal weight calibration $\Delta W^{*}$ across different projection matrices.
-Figure 3(a) shows the heatmap of $\Delta W^{*}$ derived via a Laplacian filter, revealing structured, non-random patterns that target specific channel relationships, confirming the structure-specific corrective capability of our method.
-Figure 3(b) illustrates how $\Delta W^{*}$ modulates activation responses while preserving key signal characteristics.
+The top-left panel of Figure 3 shows the heatmap of $\Delta W^{*}$ derived via a Laplacian filter, revealing structured, non-random patterns that target specific channel relationships, confirming the structure-specific corrective capability of our method.
+The top-right panel illustrates how $\Delta W^{*}$ modulates activation responses while preserving key signal characteristics.
 The smooth transitions in the corrected profile support our structure-aware calibration principle and demonstrate reduced structural distortion.
-Figure 3(c) further shows consistent error reduction across layers for various structural projection filters.
+The bottom panel further shows consistent error reduction across layers for various structural projection filters.
 DCT high-pass and learned basis filters yield superior results, especially in deeper layers where quantization errors accumulate, while random projections perform poorly, underscoring the importance of a structure-aware design.
@@
     Step 3 - Adaptive Boundary Refinement: Updates clipping boundaries $\theta$ via gradient descent while enforcing the constraint $|\text{MSE}_{x}(s^{*},\theta)-\text{MSE}_{w}(s^{*},\theta)|\leq\epsilon$.
-Table 1: Quantitative comparison of HarmoQ against SOTA quantization methods on standard SR benchmarks.
-Set5 ($\times 2$) Set14 ($\times 2$) BSD100 ($\times 2$) Urban100 ($\times 2$) Set5 ($\times 4$) Set14 ($\times 4$) BSD100 ($\times 4$) Urban100 ($\times 4$)
-Method Bit PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM PSNR/SSIM
+[…]
+[…]
 SwinIR-light 32 38.15/0.9611 33.86/0.9206 32.31/0.9012 32.76/0.9340 32.45/0.8976 28.77/0.7858 27.69/0.7406 26.48/0.7980
@@
 HarmoQ 3 38.35/0.9618 34.42/0.9250 32.36/0.9021 34.21/0.9441 31.90/0.8821 28.34/0.7700 27.41/0.7268 26.71/0.8106
+Table 1: Quantitative comparison of HarmoQ against SOTA quantization methods on standard SR benchmarks.
 Experiments
@@
 Figure 5: Visual comparison of different quantization methods on benchmark dataset.
-Table 2: Component-wise ablation analysis.
-Components Set5 ($\times 2$) Urban100 ($\times 4$)
-SRC HSO ABR Config PSNR SSIM PSNR SSIM
+[…]
+[…]
 SwinIR-Light 35.12 0.9387 23.12 0.6234
@@
 Full HarmoQ 37.98 0.9602 25.31 0.7548
+Table 2: Component-wise ablation analysis.
 Experimental Setup.
@@
 2017)) and test on five widely-used benchmarks: Set5 (Bevilacqua et al.
-2012), Set14 (Zeyde et al.
-2010), B100 (Martin et al.
-2001), Urban100 (Huang et al.
-2015), and Manga109 (Matsui et al.
+2012), Set14 (Zeyde, Elad, and Protter 2010), B100 (Martin et al.
+2001), Urban100 (Huang, Singh, and Ahuja 2015), and Manga109 (Matsui et al.
 2017).
@@
 The quantitative results validate both HarmoQ’s technical superiority and the effectiveness of coupling analysis, providing a practical pathway for efficient super-resolution model deployment.
-Table 3: Structural feature projection matrix analysis.
-Set5 ($\times 2$) Urban100 ($\times 4$)
-Projection Matrix $H$ PSNR SSIM PSNR SSIM
+[…]
+[…]
 w/o SRC 37.01 0.9521 24.45 0.6934
@@
 Random Projection 37.34 0.9548 24.78 0.7312
+Table 3: Structural feature projection matrix analysis
 Component-wise ablation.
-Our ablation study in Table 2 demonstrates that HarmoQ’s performance gains arise from complex nonlinear synergies rather than linear superposition.
+Our ablation study in Table Baseline Methods.
+demonstrates that HarmoQ’s performance gains arise from complex nonlinear synergies rather than linear superposition.
 The baseline method, which processes weight and activation quantization in isolation while ignoring their coupling, embodies the dilemma of traditional super-resolution quantization.
@@
 Structural feature projection analysis.
-Our ablation study in Table 2 explores the quest for optimal structure-aware calibration.
+Our ablation study in Table Baseline Methods.
+explores the quest for optimal structure-aware calibration.
 Laplacian filters emerge victorious (37.98/0.9602), their second-order nature perfectly capturing edge discontinuities that activation quantization disrupts.